OpenStudy (anonymous):

is 4 16 equivalent to 0.25?

OpenStudy (goldphenoix):

Correct.

OpenStudy (goldphenoix):

\[\frac{ 4 }{ 16 } = \frac{ 1 }{ 4 } = 0.25\]

OpenStudy (jkristia):

you can convert \(\frac{4}{16}\) to decimal and check if it is the same.

OpenStudy (goldphenoix):

Or you can do 4 divided by 16.

OpenStudy (jkristia):

you still need to convert 1/4 to decimal to verify